# Towards Fuzzing Zero-Knowledge Proof Circuits

### v1

@@ Abstract.
Zero-knowledge proofs (ZKPs) have evolved from a theoretical cryptographic concept into a powerful tool for implementing privacy-preserving and verifiable applications without requiring trust assumptions. Despite significant progress in the field, implementing and using ZKPs via ZKP circuits remains challenging, leading to numerous bugs that affect ZKP circuits in practice, and fuzzing remains largely unexplored as a method to detect bugs in ZKP circuits. We discuss the unique challenges of applying fuzzing to ZKP circuits, examine the oracle problem and its potential solutions, and propose techniques for input generation and test harness construction. We demonstrate that fuzzing can be effective in this domain by implementing a fuzzer for zk-regex, a cornerstone library in modern ZKP applications. In our case study, we discovered $10$ new bugs.
@@ 1. Introduction
Zero-knowledge proofs (ZKPs) are cryptographic protocols that enable a prover to convince a verifier that a given statement is true without revealing any additional information (Goldwasser et al., 1989). They are widely used in applications ranging from privacy-preserving transactions and confidential voting to secure identity verification and blockchain scalability (Ernstberger et al., 2024). To implement a ZKP, the statement must be transformed into an arithmetic circuit or constraint system. During the past decade, various domain-specific languages (DSLs) have been developed to assist developers in writing ZKP circuits (Sheybani et al., 2025).
Despite their promise, ZKPs are challenging to implement correctly. Errors in circuit design and implementation can compromise the security of a ZKP (Chaliasos et al., 2024a). Such bugs may lead to issues with completeness (failure to produce proofs for valid statements), correctness (proofs that do not accurately reflect the intended computation), or soundness (dishonest provers generating invalid proofs that are nevertheless accepted by the verifier), thereby undermining system reliability. To address these challenges, existing solutions have primarily focused on formal verification (Pailoor et al., 2023; Coglio et al., 2023a; Liu et al., 2023; Isabel et al., 2024) and static analysis (Soureshjani et al., 2023; Wen et al., 2023) to check the correctness of ZKP circuits. While these methods have significantly contributed to identifying vulnerabilities, formal verification tools often struggle with scalability when applied to complex circuits, and static analysis focuses on specific patterns while facing challenges in minimizing false positives (Chaliasos and Yu, 2025).
@@
  • We propose novel solutions to tackle the test oracle problem and the complexities of input (witness) generation for fuzzing ZKP circuits.
@@
  • We demonstrate the practical impact of fuzzing ZKP circuits by implementing a simple yet powerful fuzzer for a state-of-the-art ZKP library, which uncovered $10$ new bugs.
@@ 2. Background
A zero-knowledge proof (ZKP) is a cryptographic protocol by which a prover convinces a verifier that a given statement is true without revealing anything else beyond the truth of the statement (Goldwasser et al., 1989). A valid ZKP must satisfy three properties: (i) completeness, if the statement is true and both parties are honest, the verifier will accept the proof; (ii) soundness, if the statement is false, no cheating prover can convince the verifier to accept the proof; and (iii) zero-knowledge (ZK), the verifier learns nothing beyond the fact that the statement is true. In practice, many modern ZKPs are non-interactive and succinct. For example, zkSNARKs (zero-knowledge succinct non-interactive arguments of knowledge) (Gennaro et al., 2013) allow very short proofs and fast verification, making them practical for applications like blockchain and privacy-preserving protocols (Ernstberger et al., 2024). These protocols require the statement to be expressed as an NP-hard problem (usually as an arithmetic circuit or constraint system) (Ernstberger et al., 2023). ZKPs are widely used for private computations (e.g., shielding transaction details in Zcash (Sasson et al., 2014)) and for verifiable computations (e.g., achieving blockchain scalability with validity rollups (Chaliasos et al., 2024b)). Note that in verifiable computation, the ZK property is not required. Instead, the key advantage is that a verifier can accept a succinct proof of a potentially large or expensive computation using significantly fewer resources than re-running the entire computation. For an in-depth introduction to ZKPs, we refer the reader to Thaler et al. (2022).
ZKP Circuits. To obtain a ZKP of a statement, a developer must represent that statement in the form of a circuit. In particular, a ZKP circuit is an arithmetic representation of the computation, modeled as a collection of polynomial equations (constraints) over a finite field (Gennaro et al., 2013). These constraints define precisely which values the circuit will accept. For instance, given a function, $f$ over public and private inputs $x$ and $w$ (witness), consider the problem of proving that $y=f(x,w)$ where $y$ is claimed to be an output computed by $f$. The constraints of the circuit $C$ are constructed so that they are only satisfied by valid tuples $(x,w,y)$, i.e., by tuples where $y=f(x)$ actually holds. This allows the verifier to accept the proof. The witness $w$ can be considered a trace of the computation under consideration, and constraints are assertions that enforce the correctness of the computation. Figure 2(a) shows an example circuit, and Figures 2(b) and 2(c) illustrate an example witness and constraints for this circuit; we discuss this example in detail below.
In practice, developers use a DSL to write high-level code that describes which parts of the computation must be constrained and how. A compiler for the DSL then produces polynomial equations describing the constraints. Certain computations can be kept out of the circuit, with constraints merely validating that the external steps have been performed correctly. This approach can substantially reduce proof-generation costs. Ultimately, the circuit ensures that the verifier can check, with minimal overhead, that the prover indeed holds a valid witness for the asserted statement, all while preserving the zero-knowledge property. As noted above, the witness $w$ can be viewed as a trace of the computation. In the above example, $x$ serves as the public portion of that trace. In the case where $f(x,w)=y$, the witness $w$ contains both $x$ and $y$. Consequently, the circuit provides no zero-knowledge if $x=w$, as output should always be public.
@@
out[0] <== B*lamda*lamda - A - in1[0] - in2[0];
out[1] <== lamda * (in1[0] - out[0]) - in1[1];
@@
in1[0] = ..., in1[1] = ...
in2[0] = ..., in2[1] = ...
lamda = ...
out[0] = ..., out[1] = ...
(b) Witness
@@
Different Approaches to ZKP Development. There are three different ways that developers can use ZKPs. (i) zkDSLs such as Circom allow developers to write the circuit that implements their specifications directly. This offers tight control over efficiency but demands significant low-level expertise. (ii) Domain-specific transpilers allow developers to write higher-level statements (e.g., a regex in zk-regex (Gupta et al., 2022)) which are then transpiled into circuits (typically expressed in a zkDSL). While a good fit for specific tasks, this approach typically is not general-purpose and can introduce hidden inefficiencies and bugs if the transpiler is not correct. (iii) ZK Virtual Machines (zkVMs) like Cairo (Goldberg et al., 2021) or RISC0 (Bruestle et al., 2023) encode an entire CPU-like execution loop into some large circuits, allowing the execution of regular programs (e.g., in Rust) to be proved correct. This generality comes at the cost of performance. Notably, in all of these approaches, circuits remain a core component; what varies is who is responsible for writing them.
A critical class of bugs is soundness issues that make a circuit under-constrained, such that witnesses are not fully constrained. This situation typically violates intended functionality, and allows a malicious prover to choose a bogus witness that satisfies all the given constraints but does not correspond to a trace that the circuit can actually produce. A simple example is shown in Figure 3(b), a variant of Figure 3(a) where the result of the multiplication is unconstrained, so that the prover can convince the verifier that the circuit was used to determine bogus multiplicative relationships between numbers (e.g.,  that 2*5=100). Figure 2 shows a more realistic example. MontgomeryAdd (iden3, 2021) is used when adding two points on a Montgomery elliptic curve. Given two points (x1, y1) and (x2, y2) on the curve, a new point (xout, yout) is calculated that effectively represents their sum under elliptic-curve arithmetic. Although the computation of the circuit of Figure 2(a) is correct, this circuit is under-constrained: lambda can take any value when in2[0] == in1[0]. A malicious prover can thus convince a verifier that the addition of (x, y1) + (x + y2) can be any arbitrary values (xout, yout), even if they are not correct because those values will still satisfy the constraints that arise from the circuit, which are shown in Figure 2(c).
In contrast, an over-constrained circuit contains extraneous constraints that make the constraint system too restrictive for some legitimate inputs. These bugs lead to completeness issues. Figure 3(c) presents a variant of the safe multiplier circuit (Figure 3(a)) that suffers from a completeness bug. While the circuit works correctly when both $a$ and $b$ are 00 or both $2$, for other values, the constraints prevent the prover from constructing a valid witness, resulting in failure to generate a proof.
Table 1 outlines the different bug categories. Note that in contrast to prior work (Chaliasos et al., 2024a), we separate correctness and soundness issues, as those have different manifestations.
@@ 3. Challenges, Solutions and Open Problems of Fuzzing ZKP Programs
Applying fuzzing techniques to ZKP circuits presents unique challenges arising from their specialized nature, including uncommon bug manifestations, the separation between constraints and computation, and the heavy use of cryptographic primitives. We now discuss the main challenges in fuzzing ZKP circuits and propose potential solutions.
@@ 3.1. Key Challenges
Test Oracle Problem. A critical challenge in fuzzing is the test oracle problem. In conventional software testing, a well-defined oracle indicates whether an output is correct for a given input (Barr et al., 2014). For ZKPs, the objective is to detect completeness, correctness, and soundness bugs.
Completeness bugs are relatively straightforward to detect by verifying that valid inputs produce valid proofs, i.e., that unexpected failures are not observed. The completeness bug of Figure 3(c) could be detected by trying any inputs beyond 00 and $2$. Identifying correctness bugs requires knowledge of the expected behavior, as they manifest when an honest prover outputs a proof that, while valid in the context of the circuit, conflicts with the expected behavior. Detecting the correctness bug of Figure 3(d) would require knowing that the targeted circuit needs to perform multiplication instead of addition. Soundness bugs are significantly more challenging to detect. This is because they involve invalid proofs that the verifier erroneously accepts. Missing or insufficient constraints commonly cause such errors, as developers are working with low-level DSLs, a different programming model, and the need for optimizations. To detect the soundness bug of Figure 3(b), one would need to generate a witness that satisfies the constraints but does not match the logic of the computation.
@@
Efficiency and Performance Overheads. ZKP proof generation and verification entail computationally expensive operations. For example, proving could take up to hours for large real-world circuits. Fuzzing usually involves running large quantities of tests, but each proof can take significant time. This overhead makes fuzzing infeasible in some cases.
@@ 3.2. Proposed Solutions
(i) Mutation From Valid Witnesses: A promising approach is, to begin with a valid witness and then introduce slight modifications that make it invalid according to the oracle, thus exposing under-constrained soundness issues (Figure 5(a)). An additional improvement is to keep mutations local: if an error is introduced, adjust (i.e., patch) the remaining parts of the witness to ensure only some constraints are targeted. The intuition is that some constraints might be missing or incorrect, while others may still be valid; if the mutated value conflicts with subsequent constraints and the witness is not patched, the mismatch will cause a failure there, obscuring the bug.
(ii) Hybrid Methods: Symbolic and static analysis has been used to detect soundness bugs in ZKP circuits (Pailoor et al., 2023; Wen et al., 2023), but suffers from scalability or precision issues. We envisage a hybrid approach where fuzzing is used to find good inputs, after which the other techniques to reason about the properties of a circuit based on these inputs; see Figure 5(b). This will not provide any guarantees, but it could be a pragmatic approach to detecting bugs. The main intuition for this approach is that it will drastically reduce search space.
@@
Handling Cryptographic Primitives. A specialized oracle or white-box hooking of cryptographic components can help validate these sub-circuits. If the fuzzing target is a cryptographic function, domain-specific checks or differential testing against a reference implementation may suffice. The main difficulty is ensuring the fuzzer explores cryptographic logic deeply. For example, if a circuit takes as inputs a preimage of a hash and the hash value, we need to guide the fuzzer on how to compute the hash for a random preimage to explore the whole circuit (given that it has more functionality). The previous techniques should be applied if we want to target the hash itself.
@@ 3.3. Open Problems and Future Directions
In summary, fuzzing ZKP circuits is still in its infancy, leaving significant avenues for future research and development. Advancing this field is crucial to ensure that ZKPs remain secure and reliable as they continue to underpin an expanding range of modern applications.
@@ 4. Case Study: Fuzzing ZK-Email
To fuzz-test the circuits produced by zk-regex, we address the test oracle problem by using a reference implementation of Python’s regex module combined with differential testing. For input generation, we employ correct-by-construction and incorrect-by-construction inputs based on a specification of what constitutes a valid regex and the expected behavior of a given regex. Our campaign was conducted over one week, during which we fuzzed the library for one hour per day on a 16-core machine, continuously enhancing our fuzzer and incorporating fixes for some of the reported bugs in zk-regex. Our approach detected completeness and correctness bugs, although it did not uncover soundness bugs since we did not fuzz the witness. Despite the simplicity of our method, we revealed a total of 10 new bugs (3 correctness and 7 completeness), underscoring that even basic fuzzing campaigns can uncover significant issues in ZKP circuits. We leave as future work the implementation and evaluation of the additional techniques described in Section 3.

### v2

@@ Abstract.
Zero-knowledge proofs (ZKPs) have evolved from a theoretical cryptographic concept into a powerful tool for implementing privacy-preserving and verifiable applications without requiring trust assumptions. Despite significant progress in the field, implementing and using ZKPs via ZKP circuits remains challenging, leading to numerous bugs that affect ZKP circuits in practice, and fuzzing remains largely unexplored as a method to detect bugs in ZKP circuits. We discuss the unique challenges of applying fuzzing to ZKP circuits, examine the oracle problem and its potential solutions, and propose techniques for input generation and test harness construction. We demonstrate that fuzzing can be effective in this domain by implementing a fuzzer for zk-regex, a cornerstone library in modern ZKP applications. In our case study, we discovered $13$ new bugs that have been confirmed by the developers.
Fuzzing, Zero-Knowledge Proofs
journalyear: 2025copyright: ccconference: 34th ACM SIGSOFT International Symposium on Software Testing and Analysis; June 25–28, 2025; Trondheim, Norwaybooktitle: 34th ACM SIGSOFT International Symposium on Software Testing and Analysis (ISSTA Companion ’25), June 25–28, 2025, Trondheim, Norwaydoi: 10.1145/3713081.3731718isbn: 979-8-4007-1474-0/2025/06ccs: Security and privacy Software security engineeringccs: Security and privacy Formal methods and theory of security
@@ 1. Introduction
Zero-knowledge proofs (ZKPs) are cryptographic protocols that enable a prover to convince a verifier that a given statement is true without revealing any additional information (Goldwasser et al., 1989). They are widely used in applications ranging from privacy-preserving transactions and confidential voting to secure identity verification and blockchain scalability (Ernstberger et al., 2024b). To implement a ZKP, the statement must be transformed into an arithmetic circuit or constraint system. During the past decade, various domain-specific languages (DSLs) have been developed to assist developers in writing ZKP circuits (Sheybani et al., 2025).
Despite their promise, ZKPs are challenging to implement correctly. Errors in circuit design and implementation can compromise the security of a ZKP (Chaliasos et al., 2024a). Such bugs may lead to issues with completeness (failure to produce proofs for valid statements), correctness (proofs that do not accurately reflect the intended computation), or soundness (dishonest provers generating invalid proofs that are nevertheless accepted by the verifier), thereby undermining system reliability. To address these challenges, existing solutions have primarily focused on formal verification (Pailoor et al., 2023; Coglio et al., 2023a; Isabel et al., 2024) and static analysis (Soureshjani et al., 2023; Wen et al., 2024) to check the correctness of ZKP circuits. While these methods have contributed significantly to identifying vulnerabilities, formal verification tools often struggle with scalability when applied to complex circuits, and static analysis focuses on specific patterns while facing challenges in minimizing false positives (Chaliasos and Yu, 2025).
@@
  • We sketch initial solutions, to be further explored by future work, to tackle the test oracle problem and the complexities of input (witness) generation for fuzzing ZKP circuits.
@@
  • We demonstrate the practical impact of fuzzing ZKP circuits by implementing a simple yet powerful fuzzer for a state-of-the-art ZKP library, which uncovered $13$ new bugs.
@@ 2. Background
out[0] <== B * lambda * lambda - A - in1[0] - in2[0];
out[1] <== lambda * (in1[0] - out[0]) - in1[1];
@@
in1[0] = 1, in1[1] = 72142801481050200219322068720...
in2[0] = 4, in2[1] = 64061081989073954089143656751...
lambda = 14322771264826975277158323431...
out[0] = 27300585958536822892123621000...
out[1] = 16504465218850074729376469432...
(b) Witness (field elements)
@@
A zero-knowledge proof (ZKP) is a cryptographic protocol by which a prover convinces a verifier that a given statement is true without revealing anything else beyond the truth of the statement (Goldwasser et al., 1989). A valid ZKP must satisfy three properties: (i) completeness, if the statement is true and both parties are honest, the verifier will accept the proof; (ii) soundness, if the statement is false, no cheating prover can convince the verifier to accept the proof; and (iii) zero-knowledge (ZK), the verifier learns nothing beyond the fact that the statement is true. In practice, many modern ZKPs are non-interactive and succinct. For example, zkSNARKs (zero-knowledge succinct non-interactive arguments of knowledge) (Gennaro et al., 2013) allow very short proofs and fast verification, making them practical for applications like blockchain and privacy-preserving protocols (Ernstberger et al., 2024b). These protocols require the statement to be expressed as an NP-hard problem (usually as an arithmetic circuit or constraint system) (Ernstberger et al., 2024a). ZKPs are widely used for private computations (e.g., shielding transaction details in Zcash (Ben-Sasson et al., 2014)) and for verifiable computations (e.g., achieving blockchain scalability with validity rollups (Chaliasos et al., 2024b)). Note that in verifiable computation, the ZK property is not required. Instead, the key advantage is that a verifier can accept a succinct proof of a potentially large or expensive computation using significantly fewer resources than re-running the entire computation. For an in-depth introduction to ZKPs, we refer the reader to Thaler (2022).
ZKP Circuits. To obtain a ZKP of a statement, a developer must represent that statement in the form of a circuit. In particular, a ZKP circuit is an arithmetic representation of the computation, modeled as a collection of polynomial equations (constraints) over a finite field (Gennaro et al., 2013). These constraints define precisely which values the circuit will accept.
A circuit $C$, defines a relation $R(I,W)$, where $I$ is a public input provided by the verifier, and $W$ is a witness computed by the prover that satisfies $R$. For instance, given a function, $f$ over public input $x$, consider the problem of proving that $y=f(x)$ where $y$ is claimed to be an output computed by $f$. The constraints of the circuit $C$ are constructed so that they are only satisfied by valid tuples $(x,y)$, i.e., by tuples where $y=f(x)$ actually holds. This allows the verifier to accept the proof. The witness $W$ can be considered a trace of the computation under consideration, and constraints, i.e., the relation $R$, are assertions that enforce the correctness of the computation. In the provided example, the witness will contain both $x$, $y$.
@@
In practice, developers use a DSL to write high-level code that describes which parts of the computation must be constrained and how. A compiler for the DSL then produces polynomial equations describing the constraints. Certain computations can be kept out of the circuit, with constraints merely validating that the external steps have been performed correctly. This approach can substantially reduce proof-generation costs. Ultimately, the circuit ensures that the verifier can check, with minimal overhead, that the prover indeed holds a valid witness for the asserted statement, all while preserving the zero-knowledge property.
[…]
Different Approaches to ZKP Development. There are three different ways that developers can use ZKPs. (i) zkDSLs such as Circom allow developers to write the circuit that implements their specifications directly. This offers tight control over efficiency but demands significant low-level expertise. (ii) Domain-specific transpilers allow developers to write higher-level statements (e.g., a regex in zk-regex (Gupta et al., 2022)) which are then transpiled into circuits (typically expressed in a zkDSL). While a good fit for specific tasks, this approach is not general-purpose and can introduce hidden inefficiencies and bugs if the transpiler is not correct. (iii) ZK Virtual Machines (zkVMs) like Cairo (Goldberg et al., 2021) or RISC0 (Bruestle et al., 2023) encode an entire CPU-like execution loop into some large circuits, allowing the execution of regular programs (e.g., in Rust) to be proved correct. This generality comes at the cost of performance. Notably, in all of these approaches, circuits remain a core component; what varies is whether they are written directly by a developer, or created by upstream tooling.
@@
A critical class of bugs are soundness issues that make a circuit under-constrained, such that witnesses are not fully constrained. This situation typically violates intended functionality, and allows a malicious prover to choose a bogus witness that satisfies all the given constraints but does not correspond to a trace that the circuit can actually produce. A simple example is shown in Figure 2(b), a variant of Figure 2(a) where the result of the multiplication is unconstrained, so that the prover can convince the verifier that the circuit was used to determine bogus multiplicative relationships between numbers (e.g., that 2*5=100). Figure 1 shows a more realistic example. [^1] This example was extracted from the following report: https://veridise.com/wp-content/uploads/2023/02/VAR-circom-bigint.pdf. MontgomeryAdd (iden3, 2021) is used when adding two points on a Montgomery elliptic curve. Given two points (in1[0], in1[1]) and (in2[0], in2[1]) on the curve, a new point (out[0], out[1]) is calculated that effectively represents their sum under elliptic-curve arithmetic. Although the computation of the circuit of Figure 1(a) is correct, this circuit is under-constrained because an implicit assumption that $in2[0]-in1[0]\neq 0$. Thus, a malicious prover can convince a verifier that the addition of two points can be another value than the expected one because of the freedom to control lambda. For example, consider the scenario where all inputs are 00, then the prover can pick any value for lambda and compute the output accordingly, regardless of whether the value constitutes a correct result from the circuit. Note that the verifier will accept the proof produced because the constraints will be satisfied.
In contrast, an over-constrained circuit contains extraneous constraints that make the constraint system too restrictive for some legitimate inputs. Such bugs lead to completeness issues. Figure 2(c) presents a variant of the multiplier circuit that suffers from a completeness bug. While the circuit works correctly when both $a$ and $b$ are 00 or $2$, for other values, the constraints prevent the prover from producing a valid witness, resulting in failure to generate a proof.
Table 1 outlines the different bug categories. Note that in contrast to prior work (Chaliasos et al., 2024a), we separate correctness and soundness issues, as they manifest differently.
@@ 3. Challenges, Solutions and Open Problems of Fuzzing ZKP Programs
Applying fuzzing techniques to ZKP circuits presents unique challenges arising from their specialized nature, including uncommon bug manifestations and the separation between constraints and computation. We now discuss the main challenges in fuzzing ZKP circuits and propose potential solutions.
@@ 3.1. Key Challenges
Test Oracle Problem. A critical challenge in fuzzing is the test oracle problem. In conventional software testing, a well-defined oracle indicates whether an output is correct for a given input (Barr et al., 2015). For ZKPs, the objective is to define oracles that can detect completeness, correctness, and soundness bugs.
Completeness bugs are relatively straightforward to detect by verifying that valid inputs produce valid proofs, i.e., that unexpected failures are not observed. The completeness bug of Figure 2(c) could be detected by trying any inputs beyond 00 and $2$. Identifying correctness bugs requires knowledge of the expected behavior, as they manifest when an honest prover outputs a proof that, while valid in the context of the circuit, conflicts with the expected behavior. Detecting the correctness bug of Figure 2(d) would require knowing that the circuit under test needs to perform multiplication instead of addition. Soundness bugs are harder to detect because they involve invalid proofs that the verifier erroneously accepts. Missing or insufficient constraints commonly cause such errors, as developers are working with low-level DSLs, a different programming model, and the need for optimizations. To detect the soundness bug of Figure 2(b), one would need to generate a witness that satisfies the constraints but does not match the logic of the computation.
@@
Efficiency and Performance Overheads. ZKP proof generation and verification entail computationally expensive operations. For example, proving could take up to hours for large real-world circuits. Fuzzing usually involves running large quantities of tests, but each proof can take significant time. This overhead renders standard fuzzing techniques infeasible in some cases.
@@ 3.2. Proposed Solutions
(i) Mutation From Valid Witnesses: A promising approach is to begin with a valid witness and then introduce slight modifications that make it invalid according to the oracle, thus exposing under-constrained soundness issues (Figure 5(a)). An additional improvement is to keep mutations local: if an error is introduced, adjust (i.e., patch) the remaining parts of the witness to ensure only some constraints are targeted. The intuition is that some constraints might be missing or incorrect, while others may still be valid; if the mutated value conflicts with subsequent constraints and the witness is not patched, the mismatch will cause a failure there, obscuring the bug.
(ii) Hybrid Methods: Symbolic and static analysis has been used to detect soundness bugs in ZKP circuits (Pailoor et al., 2023; Wen et al., 2024), but suffers from scalability or precision issues. We envisage a hybrid approach where fuzzing is used to find good inputs, after which the other techniques to reason about the properties of a circuit based on these inputs; see Figure 5(b). This will not provide any guarantees, but it could be a pragmatic approach to detecting bugs. The main intuition for this approach is that it will drastically reduce search space.
@@
Handling Cryptographic Primitives. A specialized oracle or white-box hooking of cryptographic components can help validate these sub-circuits. If the fuzzing target is a cryptographic function, domain-specific checks or differential testing against a reference implementation may suffice. The main difficulty is ensuring the fuzzer explores cryptographic logic deeply. For example, if a circuit takes as inputs a preimage of a hash and the hash value, we need to guide the fuzzer on how to compute the hash for a random preimage to explore the whole circuit (given that it has more functionality).
@@ 3.3. Open Problems and Future Directions
In summary, fuzzing ZKP circuits is still in its infancy, leaving significant avenues for future research. Advancing this field is crucial to ensure that ZKPs remain secure and reliable as they continue to underpin an expanding range of modern applications.
@@ 4. Case Study: Fuzzing ZK-Email
To fuzz-test the circuits produced by zk-regex, we address the test oracle problem by using a reference implementation of Python’s regex module combined with differential testing. For input generation, we employ correct-by-construction and incorrect-by-construction inputs based on a specification of what constitutes a valid regex and the expected behavior of a given regex. Figure 6 depicts the end-to-end workflow of our fuzzer. We begin by writing a BNF specification of the fragment of regular-expression syntax supported by zk-regex. Then we employ a grammar-based fuzzer along with a pre-seeded corpus of important regexes, to get valid regexes. For each regex we then construct two types of test strings: valid inputs, which should be accepted, are obtained both by grammar expansion of the regex by employing a grammar-based fuzzer and by enumerating accepting paths in the corresponding NFA, whereas invalid inputs, which should be rejected, are generated through random mutation of valid inputs, exploration of non-accepting NFA branches, and generation from the grammar of the regex’s complement. Next, each regex is wrapped in a JSON expected by zk-regex, compiled to a Circom circuit, and executed by using valid and invalid inputs. The harness performs differential testing between Python’s re module and that of the generated circuit, and any disagreement is reported as a bug based on the used oracle that depends on if the input is valid or invalid.
Our campaign was conducted over one week, during which we fuzzed the library for one hour per day on a 16-core machine, continuously enhancing our fuzzer and incorporating fixes for some of the reported bugs in zk-regex. Our approach detected completeness and correctness bugs, although it did not uncover soundness bugs since we did not fuzz the witness. Despite the simplicity of our method, we revealed a total of 13 new bugs (5 correctness and 8 completeness), underscoring that even basic fuzzing campaigns can uncover significant issues in ZKP circuits. Notably, all $13$ bugs were confirmed by the developers of zk-regex. We leave as future work the implementation and evaluation of the additional techniques described in Section 3.